Project

General

Profile

Actions

Ops/Comm Scheduling meeting : 21/2/2017

Location : CR21 in R1 at 2pm

Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402

  • SS Magnet status
  • Hydrogen status
  • Operations
    • MOM report
    • Water chiller problem
      • The Event
    • EMR DAQ problem
    • Controls and Monitoring status
    • List of issues that need attention
  • Alarm Handler: what is the difference between "E" and "V" white alarms? Not covered in shifter training. Can they really just be ignored? AK to explain; add to shifter training
  • #1878, #1891 Trench water pump control - pressure appears stable; shifter lockout in progress
  • #1881 Water system PVs not being archived - Services/Environment/Computing spreadsheet
  • #1903, #1881 micestore disk full, also needs to be in Nagios warning checks in place but need testing (hence old data left in place)
  • #1881 File Compactor failure - shifters reported only the "previous error" message; should be in Alarm Handler. Move File Compactor output to top of RC screen? Also waiting on Services/Environment/Computing spreadsheet
  • Diffuser iris 4 not opening - not reproduced, SSU high-field test still needed
  • #1888 air pressure reading zero in diffuser GUI - fixed ad-hoc by PMH; is fix in production? C&M to close ticket when done
  • DS current - the dynamically generated Alarm Handler limits excluded the set value - AK aware
  • target1ctl dead and so no spare target control PC; target2ctl only had older software and so not publishing all data, especially laser levels - new target1ctl in difficult progress (EO)
  • Target RATS daemon crash - add "Plot From Memory" as a PV and to ALH?
  • Hall probe failure causing C&M to misunderstand Channel settings - in progress
  • MOM DC Configuration Checklist eLog tab has out-of-date info - DR: it is NOT obsolete. Add reminder to the paper DC-MOM handover forms to fill it out (AN/DC). [It sounds like it needs post-hoc updating, also to include the Tracker Helium issues, but no clear who that action should be on]
  • webcams - images not visible on micewww, probably due to micenet lockout #1399 but low priority
  • ISIS variables and status not in EPICS - restart IOC. Possibly need to get ISIS end running.
  • Neutron monitor & restart docs - MICE to take the CMS ION to Poole to be fixed ourselves. MLCR needs to know how to restart remote viewing of the temporary monitor.
  • Data Mover certificate expires 11 Feb.
  • [Q6 losing overtemperature limit from PV]
  • [Instructions for resetting Data Mover token - in progress]
AOB:
  • Estimate the full list is 3 times as long...
  • Shifters and CE training
  • BLOC and shifter handover times

Minutes

  • CC Magnet status
    The CC magnets all warmed slightly, after quench and then losing water. Leak checking the valves. Can do electrical tests - TTK Lambda to 10A.
    FC down to 60% LHe, topping up today.
    Swapped in AMI controller from SSD M1 to SSU M1. Now swapping in spare TTK Lambda from Tracker Lab.
    More noise in SSD: if can't find anything in PSU or controller, must be inside magnet and have to live with it.
    We may try the other CC magnet settings - more force but lower FC current - next ramp.
    Someone is coming out from Brookhaven to take over from Sandor on getting the magnets running.
  • MOM Report
    Have done a third of the TOF calibration and some straight-track data.
    An EMR issue keeps stopping the DAQ. Flashed firmware on to the spare board and swapped in; seems to have fixed.
  • Water Chiller
    Chiller 1 failed before weekend - contractors in today.
    The Event:
    Moving the conventional magnet circuit to use the loading-bay chiller instead of continuing to load the roof chillers. Disconnecting the Trench heat-exchanger from the roof circuit means ~300 l/min less flow round the latter and thus the pump pressure goes up past 6.5 bar. This can be mitigated by opening up the by-passes on various loops. Unfortunately one such by-pass failed and we lost water - all cryocompressors stopped.
    The Tracker compressors in the Trench melted the pipework. One may be over-cooked, is being swapped for a spare. Compressor issue? Flow or temperature switch is not a maintenance item. Need written procedure for what to do if water fails.
    Trackers probably stayed cold enough.
    The loading-bay chiller is now cooling conventional magnets - a plate is in place in the leaking valve.
    Roof chiller 2 is OK, but chiller 1 is down.
  • Controls and Monitoring status
    Run Control used to take data last week.
    Ajit thinks he may have tracked down the cause of the spurious network messages to the CArepeater. Need an IP/port list and a reboot of the machines to fix.
    Going for a static Archiver configuration.
    Most IOCs now in PRO.
AOB:
  • Estimate the full issues list is 3 times as long as mine above . Need the full issues arising list from this run (S Boyd) to prioritise/act during shutdown.
  • Shifters and CE training
    Both members of a shift crew last week were unable to close the beamstop and start on a Controlled Entry. Unacceptable: experts (e.g. those doing water failover) ARE NOT expected to know that.
    Get all shifters to go through the CE checklist during Control Room refresher training. (Paolo)
  • BLOC and shifter handover times
    If we let BLOCs act as shifter, then the BLOCs end up rostered on the same cycle as the shifters and so you have everyone swapping at the same time, losing continuity. May need to restrict doubling-up to 3-day stints or something.

Actions

  • JBoehm to get SS operating info from PHanlet (ongoing)
  • JBoehm to get/impose consensus from SC Magnets group on running of "Central CC GUI"
  • DC to collate list of issues arising from previous running
  • Status of miceraid5 disk cleanup (stands)
  • DC to ensure Health Physics have checked temporary neutron monitor (stands)
  • SBoyd - ensure list of issues arising from current running is being collated
  • SBoyd - ensure MICE-wide discussion of MOM C&M panel
  • PFranchini - Get all shifters to go through the CE checklist during Control Room refresher training
  • Henry to get gaps in BLOC rota plugged

Updated by Nebrensky, Henry over 6 years ago ยท 10 revisions